OpenStudy (dan815):

whats the flux? vector field=xi+2zj+yk; S that portion s is part of the cylinder y^ 2 +z^2=4 in the first octant bounded by x =0,x=3,y=0,z=0(assuming surface S is oriented upward)

OpenStudy (anonymous):

Really? It's been staring you in the face, @dan815 :D \[\huge (x,y,z)\rightarrow(r,\theta, z)\] \[\huge r=\sqrt{x^2+y^2}\]\[\huge \cos(\theta)=\frac{x}r\]\[\huge\sin(\theta)=\frac{y}r\] \[\huge z=z\]

OpenStudy (dan815):

what is that?

OpenStudy (anonymous):

Converting from xyz to cylindrical.

OpenStudy (dan815):

i dont get it T_T

OpenStudy (dan815):

so how do i reprsent my cylinder if its like this

OpenStudy (dan815):



OpenStudy (anonymous):

Nice drawing ^.^ You must understand, I've never heard of flux before :)

OpenStudy (dan815):

u dont need to know that just tell me how i parametrize this cylindrical coords

OpenStudy (anonymous):

^I don't know what that means... Maybe I'm not the best person for this question :)

OpenStudy (dan815):

Use the parameterization r(u, v) = <u, 2 cos v, 2 sin v> for u in [0, 3], and v in [0, π/2].

OpenStudy (dan815):

thats what i found on the internet does that make sense to you?

OpenStudy (anonymous):

Nope :) What I found is the Gauss-Divergence Theorem. ^.^

OpenStudy (dan815):

yaa this is like that

OpenStudy (anonymous):

Well, why don't you just use that? I really hate integrals with circles through them...

OpenStudy (dan815):

@wio your assistance needed plz

OpenStudy (anonymous):

If we could just use Gauss' Divergence, it just becomes equal to \[\huge \iiint\limits_V (\nabla \cdot \vec{F})dV\]

OpenStudy (dan815):

well that comes in chapter 9.15 i am still on 9.13 so it wants me to use this method right now

OpenStudy (anonymous):

What year are you in, @dan815 ?

OpenStudy (dan815):

2

OpenStudy (anonymous):

We need someone that specialises in this sort of evil Integral madness.. definitely not me... I was just using Google earlier :)

OpenStudy (dan815):

wio is like a professor

OpenStudy (dan815):

lol

OpenStudy (anonymous):

Well, there are like 4 surfaces here, right?

OpenStudy (dan815):

he always anwers these kind of questiosn for me

OpenStudy (anonymous):

If we can't do divergence theorem then we are going to do a ton more work.

OpenStudy (dan815):

i think this is what i need to find

OpenStudy (anonymous):

You need help parametrizing the surfaces?

OpenStudy (dan815):